OpenStudy (anonymous):

Let f(x) = 9x – 2 and g(x) = –x + 3. Find f(g(x)). –9x + 27 –9x + 25 –9x + 5 –9x – 2

OpenStudy (anonymous):

Replace the value of x in g(x) with 9x - 2 What are you getting??

OpenStudy (anonymous):

I don't know!

OpenStudy (anonymous):

See g(x) = -x + 3 Here if you will replace x with 9x - 2 then, g(9x - 2) = -(9x - 2) + 3 Getting or not??

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

it was B

OpenStudy (anonymous):

f(g(x))=9(-x+3)-2 f(g(x))=-9x+27-2 f(g(x))=-9x+25

OpenStudy (anonymous):

it is B SEE I HAVE SOLVEDDDD

OpenStudy (anonymous):

IS IS B

OpenStudy (anonymous):

Yeah, @waterineyes yeah.. it's not B is it...

OpenStudy (anonymous):

Ya it is B Miss Swati Ji.. Sorry I have found g(f(x)) there..

OpenStudy (anonymous):

excuseeee me

OpenStudy (anonymous):

No I am not saying to you I am saying to nitz...

OpenStudy (anonymous):

g(x)=-x+3 f(g(x))=9g(x)-2 f(g(x))=9(-x+3)-2 f(g(x))=-9x+25

OpenStudy (anonymous):

got the answer mam?????